OpenStudy (anonymous):

Let f(x,y) be xy^2 and let the constraint be g(x,y):x+4y=4 Apply the mothod of lagrange and interpret the result

OpenStudy (anonymous):

So lagrange method is: \[\Delta f = \lambda \Delta g\]

OpenStudy (anonymous):

Only those deltas are upside down (gradients)

OpenStudy (anonymous):

lol, they are called nabla

OpenStudy (anonymous):

For this problem it looks like this: \[\Delta f = <x,2xy>\]

OpenStudy (anonymous):

The gradient symbols?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Im not seeing an option for them here lol

OpenStudy (anonymous):

\nabla

OpenStudy (anonymous):

So i'll just keep using delta

OpenStudy (anonymous):

\[\nabla \]

OpenStudy (anonymous):

Magic!

OpenStudy (anonymous):

\[\nabla F=<y^2,2y x>\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and \[\lambda /\nabla g = <\lambda,4\lambda>\]

OpenStudy (anonymous):

lambda * , not lambda /

OpenStudy (anonymous):

yep, just set up equal to each other and solve for x and y

OpenStudy (anonymous):

So since those are equal we can write the following equations: \[x=\lambda\] and \[2xy=4\lambda\]

OpenStudy (anonymous):

solving the bottom equation gives:\[y=2\lambda/x\]

OpenStudy (anonymous):

from the first equation we know lambda is x

OpenStudy (anonymous):

so we know y = 2

OpenStudy (anonymous):

plugging that into our constraint equation gives:\[x+4(2)-4=0-> x=4\]

OpenStudy (anonymous):

so \[f(4,2)\] is some minimum or maximum of the function

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

plug that into function to see if it works out

OpenStudy (anonymous):

constraint function , I mean

OpenStudy (anonymous):

no, it doesnt

OpenStudy (anonymous):

then , there is mistake somewhere

OpenStudy (anonymous):

(4)+4(2)=4

OpenStudy (anonymous):

is supposed to =4 i mean

OpenStudy (anonymous):

it equals 12

OpenStudy (anonymous):

\[y^2=\lambda 1\] \[2xy=\lambda 4\] x+4y=4

OpenStudy (anonymous):

ohh, doy

OpenStudy (anonymous):

For fx i had x, not y^2

OpenStudy (anonymous):

4y^2=2xy x+4y=4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So pretty much

OpenStudy (anonymous):

after doing the lagrange multiplier its free game to do whatever you have to to find x,y?

OpenStudy (anonymous):

y=0 , x=4 x=4/3 , y=2/3 plug these two points into function, one will be max other will be min

OpenStudy (anonymous):

how are you getting 4y^2?

OpenStudy (anonymous):

\[y^2=\lambda \] while \[2xy=4 \lambda \]

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

This is really easy

OpenStudy (anonymous):

Im thinking into too much and its becoming hard haha

OpenStudy (anonymous):

Thanks for the help

OpenStudy (anonymous):

When in doubt, don't think too much

OpenStudy (zarkon):

you won't always get a/the max and min when you do Lagrange multipliers.

OpenStudy (anonymous):

saddle point?

OpenStudy (zarkon):

sort of

OpenStudy (zarkon):

on this constraint the function f has no min

OpenStudy (zarkon):

at both of your critical values \(f(x,y)\ge 0\)

OpenStudy (zarkon):

try x=-4 and y=2

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

There is another formula that is really useful for finding extrema

OpenStudy (zarkon):

the only way we are guaranteed to find the max and min is if f is continuous and the constraint forms a compact set. (then we can use the extreme value theorem)

OpenStudy (anonymous):

http://imgur.com/CAy0U

OpenStudy (zarkon):

Lagrange Multipliers is the best way to do it...you just need to be careful when checking if you really have maxs and mins

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Like here I am still trying to solve those systems

OpenStudy (anonymous):

Ive found x=4/3

OpenStudy (anonymous):

But thats it

OpenStudy (anonymous):

what happened?

OpenStudy (anonymous):

I cant figure out how wolfram gets x=4,y=0 for that system

OpenStudy (anonymous):

I see if you multiply the top system by y and subtract you get the fractional answers for x and y

OpenStudy (anonymous):

4y^2=2xy well , this works when y=0 0=0 assuming y=0 x+4(0)=4 x=4

OpenStudy (anonymous):

Hmm. ok

OpenStudy (anonymous):

So interpreting these results, would it be correct to say: The function f(x,y) has a minimum or maximum at f(4,0) or f(4/3,2/3)